A Carnot engine absorbs heat from a source at a higher temperature of 800 Kelvin

and the efficiency of the Carnot engine is 50%.

To what value the higher temperature should be increased for efficiency to increase

to 80%, if the lower temperature is kept constant.


1
Expert's answer
2022-03-07T14:56:03-0500


"\eta" "=\frac{t_2-t_1}{t_2}\%" or "(1-\frac{T_2}{T_1})"


Where "t_2" is the source temperature

"t_1=" intake temperature


"\frac{50}{100}=\frac{800-t_1}{800}\implies\>t_1=400K"


Where the efficiency is "80\%"


"\frac{80}{100}=\frac{t_2-400}{t_2}"


"t_2=2000K"